Below is a spreadsheet for a situation in which a day trader seeks to maximize expected return while minimizing risk for a portfolio consisting of two assets.

A

B

C

D

1

Day Trading Portfolio

2

3

Probabilities and Outcomes:

Probability

X Index Fund

Y Index Fund

4

0.25

$5

$6

5

0.36

6

0.19

$3

7

0.04

$2

8

0.05

$1

$(4)

9

0.03

$3

10

0.06

$(2)

11

0.01

$(2)

$5

12

0.01

$(3)

$1

38) What is the variance of the X index fund?

A) 8.80

B) 22.78

C) 4.69

D) 16.40

Answer: C

-------------------

39) What is the standard deviation of the X index fund?

A) 5.13

B) 2.17

C) 4.77

D) 2.97

Answer: B

-------------------

40) What is the variance of the Y index fund?

A) 22.78

B) 4.69

C) 18.66

D) 8.80

Answer: D

-------------------

41) What is the standard deviation of the Y index fund?

A) 5.17

B) 4.77

C) 2.17

D) 2.97

Answer: D

-------------------

42) What is the covariance between the two assets?

A) 4.64

B) 4.69

C) 4.77

D) 4.81

Answer: A

How we got the results ??
Answer #1

38)

Var(X)=E(X^{2})-E(X)^{2}

E(X)=\sum xp(x)=5(0.25)+3(0.19)+2(0.04)+1(0.05)+(-2)(0.01)+(-3)(0.01)=1.9

E(X^{2})=\sum x^{2}p(x)=5^{2}(0.25)+3^{2}(0.19)+2^{2}(0.04)+1^{2}(0.05)+(-2)^{2}(0.01)+(-3)^{2}(0.01)=8.3

So, Var(X)=8.3-1.9^{2}=4.69

Hence option C.

39)

Standard Deviation X = Var(X)0.5 = 4.690.5 = 2.17

Hence option B.

40)

Var(Y)=E(Y^{2})-E(Y)^{2}

E(Y)=\sum yp(y)=6(0.25)+(-4)(0.05)+3(0.03)+(-2)(0.06)+(5)(0.01)+(1)(0.01)=1.33

E(Y^{2})=\sum y^{2}p(y)=6^{2}(0.25)+(-4)^{2}(0.05)+3^{2}(0.03)+(-2)^{2}(0.06)+(5)^{2}(0.01)+(1)^{2}(0.01)=10.57

So, Var(Y)=10.57-1.33^{2}=8.80

Hence option D.

41)

Standard Deviation Y = Var(Y)0.5 = 8.800.5 = 2.97

Hence option D.

42)

Cov(X,Y)=E(XY)-E(X)E(Y)

E(XY)=\sum \sum xyp(x,y)=30(0.25)+(-4)(0.05)+(-10)(0.01)+(-3)(0.01)=7.17

So, Cov(X,Y)=7.17-(1.9)(1.33)=4.64

Hence option A.